OpenStudy (anonymous):

these are really hard for me.. Can someone help? The vertex of y = 9 - 8x - x2 is (-4, 25) True or False What is the vertex for the function? y = x2 - 4x - 3 Anwsers: (-1, 2) (0, -3) (1, -6) (2, -7)

OpenStudy (anonymous):

I am assuming that when you say x2 you mean \[x^2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(2,-7)

OpenStudy (anonymous):

Can you help me with the previous ones i posted?

OpenStudy (anonymous):

To do this, we find the axis of symmetry, which (in standard form of ax^2 + bx + c) is x = -b/a which in this case is -8, then we plug that into x to get the y-value of the vertex, 9, so the vertex is (-8, 9)

OpenStudy (anonymous):

so false

OpenStudy (anonymous):

can you help me with these two? What are the values of a, b, and c in the function? List them in order with commas, such as 1,-2,3. y = -2x2 + 5x - 4 What is the vertex for the function? (You will want to find the axis of symmetry first.) Write as an ordered pair, leaving no spaces, such as (1,-4). y = -3x2 + 6x

OpenStudy (anonymous):

look at the standard form i posted above.

OpenStudy (anonymous):

i dnt get it

OpenStudy (anonymous):

\[y = ax^2 + bx + c\]

OpenStudy (anonymous):

??

OpenStudy (anonymous):

now you can do the first one, then look how I found the axis of symmetry above, -b/a, in this case on the second question -6/-3 = 2

OpenStudy (anonymous):

Ok, What about the first one?

OpenStudy (anonymous):

\[y = ax^2 + bx + c \] this is the standard form of a quadratic equation, quadratic means that the highest exponent is 2 such as \[x^2\]

OpenStudy (anonymous):

OK,

OpenStudy (anonymous):

find the coefficients of x^2 and x in their respective places relative to a and b, c is the constant number on the end (-4)

OpenStudy (anonymous):

i got it :)

OpenStudy (anonymous):

One more? Thanks so much for your help! Which coefficient in the function tells us that the parabola for the function will point downward? y = -2x2 + 5x - 4

OpenStudy (anonymous):

Anwsers: -2 5 -4

OpenStudy (anonymous):

the direction of the parabola depends on whether the coefficient of the x^2 term is positive or negative. If it is positive such as 5x^2 it will point up, but -3x^2 would point down. In this case -2 is the coefficient which means it points downward

OpenStudy (anonymous):

ohhh kk!!

OpenStudy (anonymous):

HOW DO I FIGURE THIS OUT \[6\sqrt{2} - 3\sqrt{2}\]